# Heterogeneous window transformer for image denoising

### v1

@@ Abstract
Deep networks can usually depend on extracting more structural information to improve denoising results. However, they may ignore correlation between pixels from an image to pursue better denoising performance. Window transformer can use long- and short-distance modeling to interact pixels to address mentioned problem. To make a tradeoff between distance modeling and denoising time, we propose a heterogeneous window transformer (HWformer) for image denoising. HWformer first designs heterogeneous global windows to capture global context information for improving denoising effects. To build a bridge between long and short-distance modeling, global windows are horizontally and vertically shifted to facilitate diversified information without increasing denoising time. To prevent the information loss phenomenon of independent patches, sparse idea is guided a feed-forward network to extract local information of neighboring patches. The proposed HWformer only takes 30% of popular Restormer in terms of denoising time.
@@ I Introduction
Convolutional neural networks (CNNs) can obtain strong learning abilities via stacking simple components, i.e., convolutional layers and activation functions to overcome mentioned challenges for video and image applications [5, 6], especially image denoising [7]. Denoising CNN (DnCNN) uses convolutional layers, batch normalization techniques, rectified linear unit (ReLU) and batch normalization techniques to make a tradeoff between denoising performance and efficiency [7]. To promote denoising effects, residual dense network (RDN) integrated hierarchical information to facilitate more richer detailed information for recovering clean images [8]. Moreover, asymmetric CNN (ACNet)[9] embed asymmetric ideas into CNN to enhance local features to restore more details for image resotration. Due to small kernels, CNNs can only obtain local information, which may limit denoising performance. Transformer uses pixel relations to capture global information to overcome drawback of CNNs for image denoising [10]. Image process Transformer (IPT) utilizes a Transformer containing a self-attention mechanism and a feed-forward network to interact global pixels for promoting denoising effects [10]. Alternatively, SwinIR restricts effect area of a self-attention mechanism to a local window to reduce computational cost for image denoising [11].
Although the window Transformer can rely on a short-distance modeling to reduce denoising time, local windows can limit interactions of contexts. Also, existing transformer based long-distance modeling will increase denoising time. To build a bridge between distance modeling and denoising time, we present a heterogeneous window transformer (HWformer) in image denoising. HWformer first designs heterogeneous global windows to try best to guarantee interactions of self-attention mechanisms to obtain more global context information for promoting performance of image denoising. To break the limits of interactions of long and short-distance modeling, global windows are horizontally and vertically shifted to facilitate diversified information without increasing denoising time. To prevent native effects of independent patches, sparse idea is first embedded into a feed-forward network to extract more local information of neighboring patches. Also, our HWformer only takes 30% of popular Restormer in terms of denoising time.
@@
1.Heterogeneous global windows with different sizes are designed to facilitate richer global context information to overcome drawback of short-distance modeling.
@@ II Related work
Although convolutional neural networks have powerful feature extraction abilities, they are still faced with challenges of data scarcity and imbalance for image denoising. In terms of data scarcity, data augmentation techniques are good choice for image denoising [12]. Due to strong generative abilities, generative adversarial networks (GANs) are used for data augmentation to improve performance of image denoising [13]. To increase the number of training data, Wasserstein is embedded into a GAN to improve denoising effect [14]. six paper. Yang et al.[13] used optimal transport idea to enhance a GAN to enhance data for promoting denoising effect. To address real noisy image denoising, a two-step denoising method is presented [15]. That is, the first step maps given images as raw images. The second step uses a GAN to estimate noise distribution on a large scale collected images, where obtained noise distribution can be used to train a denoising model. Alternatively, Hong et al. [16] utilized conditional generative adversarial network to learn noise distribution from given noisy images to achieve a image blind denoising. Fuentes et al. [17] fused a regularization term into a GAN to learn noise distribution and used structure preserving loss to improve denoising results. In terms of data imbalance, normalization techniques [18] are effective tools to address nonuniform data distribution in image denoising. For instance, Zhang et al. applied batch normalization techniques, a residual learning operation to act a convolutional network to achieve an efficient denoising model, where batch normalization technique is used to normalize obtained structural information to make a tradeoff between training efficiency and denoising performance. To address non-uniformly distributed data affected by constrained-resource hardware platform, Tian et al. [19] used batch renormalization technique to normalize whole sample rather than normalizing batch of batch normalization to improve denoising effect. To prevent overfitting, adaptive instance normalization is gathered into a CNN to establish a denoiser [20]. Mentioned illustrations can use structural information to improve performance of image denoising. However, pixel relations can improve effects of image denoising. Thus, we combine structural relation and pixel relation to improve denosing results in this paper.
@@ III-C Global-window Transformer enhancement block
To break the limits of short-distance modeling, GTEBlock uses global window rather than local window to extract global information. That is, GTEBlock cuts into windows of $96\times 96$ rather than that of $48\times 48$ [10] in the self-attention to enlarge receptive field for capturing more global information, which can also break the limits of global information loss of short-distance modeling. That is, obtained 2D features input into three independent convolutional layers rather than fully connected layers to obtain query (Q), key (K) and value (V), where used convolutional layers can reduce parameters. Then, GTEBlock cuts respectively Q, K and V to non-overlapping patches of $p\times p$ to reduce computational costs and they are flattened as vectors. Next, to make stable training, obtained vectors are normalized [23] as inputs of a multi-head self-attention mechanism to achieve global interactions of pixels, where the multi-head self-attention mechanism can be obtained at Ref. [11]. Finally, a combination of two convolutional layers and ReLU is rather than 2-layer feedforward network to reduce the number of parameters. To prevent long-term dependency problem, two residual learning operations are applied in each GTEBlock as shown in Fig.1. Additionally, its pseudocode can be shown in Algorithm 2.
@@ IV-B Ablation study
TABLE III: Average PSNR(dB) of different gray image denoising methods on Set12[7] with different noise levels of 15, 25 and 50.
Images C.man House Peppers Starfish Monarch Airplane Parrot Lena Barbara Boat Man Couple Average
@@
BM3D 31.91 34.93 32.69 31.14 31.85 31.07 31.37 34.26 33.10 32.13 31.92 32.10 32.37
TNRD 32.19 34.53 33.04 31.75 32.56 31.46 31.63 34.24 32.13 32.14 32.23 32.11 32.50
DnCNN 32.61 34.97 33.30 32.20 33.09 31.70 31.83 34.62 32.64 32.42 32.46 32.47 32.86
FFDNet 32.43 35.07 33.25 31.99 32.66 31.57 31.81 34.62 32.54 32.38 32.41 32.46 32.77
FOCNet 32.71 35.44 33.41 32.40 33.29 31.82 31.98 34.85 33.09 32.62 32.56 32.64 33.07
RDDCNN 32.61 35.01 33.31 32.13 33.13 31.67 31.93 34.57 32.62 32.42 32.38 32.46 32.85
DGAL 32.73 35.83 33.50 32.57 33.56 31.96 32.08 34.89 33.81 32.68 32.61 32.73 33.25
CTNet 32.82 35.86 33.69 32.65 33.53 32.07 32.21 34.90 33.87 32.75 32.61 32.77 33.31
SwinIR 32.93 36.00 33.72 32.59 33.66 32.12 32.22 34.97 33.97 32.80 32.68 32.84 33.38
HWformer (Ours) 32.93 36.21 33.80 32.78 33.68 32.06 32.25 34.96 34.08 32.86 32.65 32.86 33.43
@@
BM3D 29.45 32.85 30.16 28.56 29.25 28.42 28.93 32.07 30.71 29.90 29.61 29.71 29.97
TNRD 29.72 32.53 30.57 29.02 29.85 28.88 29.18 32.00 29.41 29.91 29.87 29.71 30.06
DnCNN 30.18 33.06 30.87 29.41 30.28 29.13 29.43 32.44 30.00 30.21 30.10 30.12 30.43
FFDNet 30.10 33.28 30.93 29.32 30.08 29.04 29.44 32.57 30.01 30.25 30.11 30.20 30.44
N3Net 30.08 33.25 30.90 29.55 30.45 29.02 29.45 32.59 30.22 30.26 30.12 30.12 30.55
FOCNet 30.35 33.63 31.00 29.75 30.49 29.26 29.58 32.83 30.74 30.46 30.22 30.40 30.73
RDDCNN 30.20 33.13 30.82 29.38 30.36 29.05 29.53 32.40 30.03 30.19 30.05 30.10 30.44
DGAL 30.36 33.88 31.18 29.99 30.77 29.40 29.65 32.88 31.56 30.51 30.28 30.45 30.91
CTNet 30.40 33.86 31.33 30.03 30.68 29.50 29.73 32.87 31.62 30.54 30.27 30.49 30.94
SwinIR 30.53 33.99 31.35 29.98 30.89 29.55 29.77 32.94 31.70 30.63 30.33 30.59 31.02
HWformer (Ours) 30.51 34.21 31.41 30.23 30.98 29.48 29.83 32.96 31.83 30.68 30.31 30.61 31.09
@@
BM3D 26.13 29.69 26.68 25.04 25.82 25.10 25.90 29.05 27.22 26.78 26.81 26.46 26.72
TNRD 26.62 29.48 27.10 25.42 26.31 25.59 26.16 28.93 25.70 26.94 26.98 26.50 26.81
DnCNN 27.03 30.00 27.32 25.70 26.78 25.87 26.48 29.39 26.22 27.20 27.24 26.90 27.18
FFDNet 27.05 30.37 27.54 25.75 26.81 25.89 26.57 29.66 26.45 27.33 27.29 27.08 27.32
N3Net 27.14 30.50 27.58 26.00 27.03 25.75 26.50 29.76 27.01 27.32 27.33 27.04 27.43
FOCNet 27.36 30.91 27.57 26.19 27.10 26.06 26.75 29.98 27.60 27.53 27.42 27.39 27.68
RDDCNN 27.16 30.21 27.38 25.72 26.84 25.88 26.53 29.32 26.36 27.23 27.22 26.88 27.23
DGAL 27.30 31.09 27.78 26.37 27.16 26.16 26.76 29.91 28.27 27.50 27.44 27.36 27.76
CTNet 27.47 30.98 27.92 26.45 27.14 26.28 26.70 29.89 28.29 27.52 27.41 27.37 27.79
SwinIR 27.47 31.25 27.91 26.45 27.24 26.25 26.95 30.09 28.39 27.65 27.50 27.54 27.89
HWformer (Ours) 27.54 31.44 27.96 26.76 27.43 26.22 26.92 30.08 28.56 27.72 27.51 27.55 27.97
@@
It is known that existing short-distance modeling will lose global information. To overcome the drawback, we design a heterogeneous architecture containing two GTEBlocks and one TDEBlock with proposed sparse technique to build bridge for interacting between short- and long-distance modeling in image denoising. Thus, in this subsection, we give principle and effectiveness of key techniques, i.e., two GTEBlocks, TDEBlock and sparse technique in the HWformer as follows.
@@
Global-window Transformer enhancement block: Taking into inferiority of short-distance modeling, we use windows of $96\times 96$ rather than that of $48\times 48$ as an input of GTEBlock to capture more global information. ID 7 of GTEBlock has obtained an improvement of 0.08dB in terms of PSNR than that ID 1 for image denoising in TABLE I, which shows superiority of enlarging windows in our HWformer for image denoising. Also, improvement of 0.1dB is not easy for image denoising. Thus, it has a reference value for improving Transformer for image denoising. Because Transformer mechanism can only depend on relations of pixels to extract salient information [10], three fully connected layers in the Transformer mechanism depend on sizes of given patches, which will result in high complexities when given patches is big. To overcome this issue, we use three convolutional layers with sizes of fixed kernels rather than three fully connected layers to eliminate native effects of three fully connected layers in terms of the number of parameters for image denoising. Its effectiveness can be verified as follows. We assume that parameters of each fully-connected layer are $C\times p^{2}\times C\times p^{2}$, where $C$ is channel number, $p$ is size of given patches ($p=6$ in the GTEBlock). According to that, we see that parameters of each fully-connected layer are strong relation with $p$. However, parameters of each convolutional layer are $C\times C\times k\times k$ , where $k$ denotes kernel size ($k=3$). According to parameter computation method, we can see that, the number of parameters is not affected by given patch sizes. Also, the number of parameters from each convolutional layer only takes 1/144 that of each fully connected layer, which can reduce computational costs. Besides, its performance does not any varying via ID 5 of GTEBlock with fully connected layers (FCL) and ID 6 of GTEBlock as shown in TABLE I. We can see that GTEBlock with three fully connected layers has obtained same PSNR as GTEBlock with three convolutional layers on Urban100 [24] for image denoising with noise level of 15.
@@ IV-C Experimental results
We use quantitative and qualitative analysis to test denoising performance of our HWformer. Quantitative analysis includes synthetic and real noisy image denoising, comparisons of complexity and evaluation of denoising time. Qualitative analysis is used to evaluate visual effects of predicted denoising images. For synthetic noisy image denoising, we use Block-matching and 3D filtering (BM3D) [38], trainable nonlinear reaction diffusion (TNRD) [39], targeted image denoising (TID) [40], a denoising CNN (DnCNN) [7], image restoration CNN (IRCNN) [41], fast and flexible denoising network (FFDNet) [42], neural nearest neighbors networks (N3Net) [43], fractional optimal control network (FOCNet) [44], DRUNet [45], DAGL [46], robust deformed denoising CNN (RDDCNN) [47], cross Transformer denoising CNN (CTNet) [30], CSformer [48], image restoration using swin Transformer (SwinIR) [11], Restoration Transformer (Restormer) [49], image processing transformer (IPT) [10] and encoder-decoder-based transformer (EDT) [50] on Urban100 [24], BSD68 [31] and Set12 [7] for gray synthetic noisy image denoising, on CBSD68 [31], Kodak24 [33], SPARCS [34], McMaster [32] and Urban100 [24] for color synthetic noisy image denoising. To comprehensively validate the proposed method in image denoising, peak signal-to-noise ratio (PSNR)[51], structural similarity index (SSIM)[51], feature similarity index measure (FSIM)[52], learned perceptual image patch similarity (LPIPS) [53], peak signal-to-blur ratio (PSBR)[54] and color difference [55] are used as metric to conduct experiments. For gray image denoising, experiments of different methods for images denoising with single class and multiple classes are conducted. As shown in TABLE III, our HWformer method almost has obtained the best result for single image denoising with twelve classes when noise level is 15, 25 and 50, respectively. For instance, our HWformer exceeds 0.31dB than that of the second SwinIR for a Starfish image denoising with noise level of 50 in TABLE III. As shown in TABLE IV, we can see that our HWformer has nearly obtained the highest PSNR on BSD68 and Urban 100 for image denoising with muplti classes when noise level is 15, 25 and 50, respectively. For instance, our HWformer has an imporvement of 0.15dB than that of the second Restormer on Urban 100 for noise level of 25 in TABLE IV. According to Mentioned illusrations, it is known that our HWformer is effective for gray image denoising.
@@
For real noisy image denoising, we compare with popular methods, i.e., BM3D [38], TID [40], DnCNN [7], RIDNet [56], VDN [57], AINDNet [20], SADNet [58], RDDCNN [47], DANet [59], DeamNet [60], CycleISP [61], DAGL [46] and MPRNet [62] on CC [35] and SIDD [29]. Specifically, we use given reference images (Ground truth) from chosen real noisy image datasets and obtained clean images from different noisy models to compute PSNR [51] values to test denoising performance of our HWformer. As shown in TABLE X, our method have achieves the best performance besides one settings of Nikon D800 ISO = 1600 on CC for real noisy image denoising. For instance, our method has an improvement of 3.38dB than that of the second CBM3D in terms of PSNR values of the third image with the setting of Nikon D600 ISO = 3200 for real noisy image denoising in TABLE X. As TABLE XI, we can see that the proposed HWformer is competitive in contrast with other popular methods, i.e., ADGL and MPRNet on SIDD for real noisy image denoising.Also, red and blue lines denote the best and second results in TABLEs III-XI.
@@
Qualitative analysis: We choose an area of denoising images from different method to amplify it as observation area. If observation area is clearer, its corresponding method is more effective for image denoising. Also, we choose DnCNN, SwinIR, Restormer, DRUNet, DGAL and our HWformer on BSD68 and Urban100 for noise levels of 50 to obtain gray clean images. As shown in Figs. 2 and 3, we can see that our HWformer has obtained clearer areas than that of other comparative methods. For comparisons of color visual images, we choose DnCNN, SwinIR, Restormer, DRUNet, EDT-B and our HWformer on McMaster and Urban100 for noise levels of and 50 to obtain color clean images. As shown in Figs. 4 and 5, we can see that our HWformer can obtain clearer detailed information than that of other methods. For real image denoising, we can observe our HWformer is capable of restoring more details and texture information as shown in Fig. 6. Thus, our HWformer is effective for qualitative analysis.

### v2

@@ Abstract
Deep networks can usually depend on extracting more structural information to improve denoising results. However, they may ignore correlation between pixels from an image to pursue better denoising performance. Window Transformer can use long- and short-distance modeling to interact pixels to address mentioned problem. To make a tradeoff between distance modeling and denoising time, we propose a heterogeneous window Transformer (HWformer) for image denoising. HWformer first designs heterogeneous global windows to capture global context information for improving denoising effects. To build a bridge between long and short-distance modeling, global windows are horizontally and vertically shifted to facilitate diversified information without increasing denoising time. To prevent the information loss phenomenon of independent patches, sparse idea is guided a feed-forward network to extract local information of neighboring patches. The proposed HWformer only takes 30% of popular Restormer in terms of denoising time. Its codes can be obtained at https://github.com/hellloxiaotian/HWformer.
@@ I Introduction
Convolutional neural networks (CNNs) can obtain strong learning abilities via stacking simple components, i.e., convolutional layers and activation functions to overcome mentioned challenges for video and image applications [5, 6], especially image denoising [7]. Denoising CNN (DnCNN) uses convolutional layers, batch normalization techniques, rectified linear unit (ReLU) and batch normalization techniques to make a tradeoff between denoising performance and efficiency [7]. To promote denoising effects, residual dense network (RDN) integrated hierarchical information to facilitate richer detailed information for recovering clean images [8]. Moreover, asymmetric CNN (ACNet)[9] embedded asymmetric ideas into a CNN to enhance local features to restore more details for image resotration. Due to small kernels, CNNs can only obtain local information, which may limit denoising performance. Transformer uses pixel relations to capture global information to overcome drawbacks of CNNs for image denoising [10]. Image process Transformer (IPT) utilizes a Transformer containing a self-attention mechanism and a feed-forward network to interact global pixels for promoting denoising effects [10]. Alternatively, SwinIR restricts effect area of a self-attention mechanism to a local window to reduce computational cost for image denoising [11].
Although the window Transformer can rely on a short-distance modeling to reduce denoising time, local windows can limit interactions of contexts. Also, existing Transformer based long-distance modeling will increase denoising time. To build a bridge between distance modeling and denoising time, we present a heterogeneous window Transformer (HWformer) in image denoising. HWformer first designs heterogeneous global windows to try best to guarantee interactions of self-attention mechanisms to obtain more global context information for promoting performance of image denoising. To break the limitation of interactions of long and short-distance modeling, global windows are horizontally and vertically shifted to facilitate diversified information without increasing denoising time. To prevent native effects of independent patches, sparse idea is first embedded into a feed-forward network to extract more local information of neighboring patches. Also, our HWformer only takes 30% of popular Restormer in terms of denoising time.
@@
1. Heterogeneous global windows with different sizes are designed to facilitate richer global context information to overcome drawback of short-distance modeling.
@@ II Related work
Although convolutional neural networks have powerful feature extraction abilities, they are still faced with challenges of data scarcity and imbalance for image denoising. In terms of data scarcity, data augmentation techniques are good choice for image denoising [12]. Due to strong generative abilities, generative adversarial networks (GANs) are used for data augmentation to improve performance of image denoising [13]. To increase the number of training data, Wasserstein is embedded into a GAN to improve denoising effect [14]. Yang et al.[13] used optimal transport idea to enhance a GAN to enhance data for promoting denoising effect. To address real noisy image denoising, a two-step denoising method is presented [15]. That is, the first step maps given images as raw images. The second step uses a GAN to estimate noise distribution on a large scale collected images, where obtained noise distribution can be used to train a denoising model. Alternatively, Hong et al. [16] utilized conditional generative adversarial network to learn noise distribution from given noisy images to achieve a image blind denoising. Fuentes et al. [17] fused a regularization term into a GAN to learn noise distribution and used structure preserving loss to improve denoising results. In terms of data imbalance, normalization techniques [18] are effective tools to address nonuniform data distribution in image denoising. For instance, Zhang et al. applied batch normalization techniques, a residual learning operation to act a convolutional network to achieve an efficient denoising model, where batch normalization technique is used to normalize obtained structural information to make a tradeoff between training efficiency and denoising performance. To address non-uniformly distributed data affected by constrained-resource hardware platform, Tian et al. [19] used batch renormalization technique to normalize whole sample rather than normalizing batch of batch normalization to improve denoising effect. To prevent overfitting, adaptive instance normalization is gathered into a CNN to establish a denoiser [20]. Mentioned illustrations can use structural information to improve performance of image denoising. However, pixel relations can improve effects of image denoising. Thus, we combine structural relation and pixel relation to improve denosing results in this paper.
@@ III-C Global-window Transformer enhancement block
To break the limits of short-distance modeling, GTEBlock uses global window rather than local window to extract global information. That is, GTEBlock cuts into windows of $96\times 96$ rather than that of $48\times 48$ [10] in the self-attention to enlarge receptive field for capturing more global information, which can also break the limits of global information loss of short-distance modeling. That is, obtained 2D features input into three independent convolutional layers rather than fully connected layers to obtain query (Q), key (K) and value (V), where used convolutional layers can reduce parameters. Then, GTEBlock cuts respectively Q, K and V to non-overlapping patches of $p\times p$ to reduce computational costs and they are flattened as vectors. Next, to make training stable, obtained vectors are normalized [23] as inputs of a multi-head self-attention mechanism to achieve global interactions of pixels, where the multi-head self-attention mechanism can be obtained at Ref. [11]. Finally, a combination of two convolutional layers and ReLU is rather than 2-layer feedforward network to reduce the number of parameters. To prevent long-term dependency problem, two residual learning operations are applied in each GTEBlock as shown in Fig.1. Additionally, its pseudocode can be shown in Algorithm 2.
@@ IV-B Ablation study
TABLE III: Average PSNR(dB) of eleven gray image denoising methods on Set12[7] with different noise levels of 15, 25 and 50.
Images C.man House Peppers Starfish Monarch Airplane Parrot Barbara Boat Man Couple Average
@@
BM3D 31.91 34.93 32.69 31.14 31.85 31.07 31.37 33.10 32.13 31.92 32.10 32.20
TNRD 32.19 34.53 33.04 31.75 32.56 31.46 31.63 32.13 32.14 32.23 32.11 32.34
DnCNN 32.61 34.97 33.30 32.20 33.09 31.70 31.83 32.64 32.42 32.46 32.47 32.69
FFDNet 32.43 35.07 33.25 31.99 32.66 31.57 31.81 32.54 32.38 32.41 32.46 32.60
FOCNet 32.71 35.44 33.41 32.40 33.29 31.82 31.98 33.09 32.62 32.56 32.64 32.91
RDDCNN 32.61 35.01 33.31 32.13 33.13 31.67 31.93 32.62 32.42 32.38 32.46 32.70
DGAL 32.73 35.83 33.50 32.57 33.56 31.96 32.08 33.81 32.68 32.61 32.73 33.10
CTNet 32.82 35.86 33.69 32.65 33.53 32.07 32.21 33.87 32.75 32.61 32.77 33.20
SwinIR 32.93 36.00 33.72 32.59 33.66 32.12 32.22 33.97 32.80 32.68 32.84 33.23
HWformer (Ours) 32.93 36.21 33.80 32.78 33.68 32.06 32.25 34.08 32.86 32.65 32.86 33.29
@@
BM3D 29.45 32.85 30.16 28.56 29.25 28.42 28.93 30.71 29.90 29.61 29.71 29.78
TNRD 29.72 32.53 30.57 29.02 29.85 28.88 29.18 29.41 29.91 29.87 29.71 29.88
DnCNN 30.18 33.06 30.87 29.41 30.28 29.13 29.43 30.00 30.21 30.10 30.12 30.25
FFDNet 30.10 33.28 30.93 29.32 30.08 29.04 29.44 30.01 30.25 30.11 30.20 30.31
N3Net 30.08 33.25 30.90 29.55 30.45 29.02 29.45 30.22 30.26 30.12 30.12 30.31
FOCNet 30.35 33.63 31.00 29.75 30.49 29.26 29.58 30.74 30.46 30.22 30.40 30.53
RDDCNN 30.20 33.13 30.82 29.38 30.36 29.05 29.53 30.03 30.19 30.05 30.10 30.26
DGAL 30.36 33.88 31.18 29.99 30.77 29.40 29.65 31.56 30.51 30.28 30.45 30.73
CTNet 30.40 33.86 31.33 30.03 30.68 29.50 29.73 31.62 30.54 30.27 30.49 30.77
SwinIR 30.53 33.99 31.35 29.98 30.89 29.55 29.77 31.70 30.63 30.33 30.59 31.85
HWformer (Ours) 30.51 34.21 31.41 30.23 30.98 29.48 29.83 31.83 30.68 30.31 30.61 30.92
@@
BM3D 26.13 29.69 26.68 25.04 25.82 25.10 25.90 27.22 26.78 26.81 26.46 26.51
TNRD 26.62 29.48 27.10 25.42 26.31 25.59 26.16 25.70 26.94 26.98 26.50 26.62
DnCNN 27.03 30.00 27.32 25.70 26.78 25.87 26.48 26.22 27.20 27.24 26.90 26.98
FFDNet 27.05 30.37 27.54 25.75 26.81 25.89 26.57 26.45 27.33 27.29 27.08 27.10
N3Net 27.14 30.50 27.58 26.00 27.03 25.75 26.50 27.01 27.32 27.33 27.04 27.2
FOCNet 27.36 30.91 27.57 26.19 27.10 26.06 26.75 27.60 27.53 27.42 27.39 27.44
RDDCNN 27.16 30.21 27.38 25.72 26.84 25.88 26.53 26.36 27.23 27.22 26.88 27.04
DGAL 27.30 31.09 27.78 26.37 27.16 26.16 26.76 28.27 27.50 27.44 27.36 27.56
CTNet 27.47 30.98 27.92 26.45 27.14 26.28 26.70 28.29 27.52 27.41 27.37 27.59
SwinIR 27.47 31.25 27.91 26.45 27.24 26.25 26.95 28.39 27.65 27.50 27.54 27.69
HWformer (Ours) 27.54 31.44 27.96 26.76 27.43 26.22 26.92 28.56 27.72 27.51 27.55 27.78
@@
It is known that existing short-distance modeling will lose global information to reduce effect of image processing. To overcome the drawback, we design a heterogeneous architecture containing two GTEBlocks and one TDEBlock with proposed sparse technique to build bridge for interacting between short- and long-distance modeling in image denoising. Thus, in this subsection, we give principle and effectiveness of key techniques, i.e., two GTEBlocks, TDEBlock and sparse technique in the HWformer as follows.
@@
Global-window Transformer enhancement block: Taking into inferiority of short-distance modeling account, we use windows of $96\times 96$ rather than that of $48\times 48$ as an input of GTEBlock to capture more global information. ID 7 of GTEBlock has obtained an improvement of 0.08dB in terms of PSNR than that ID 1 for image denoising in TABLE I, which shows superiority of enlarging windows in our HWformer for image denoising. Also, improvement of 0.1dB is not easy for image denoising. Thus, it has a reference value for improving Transformer for image denoising. Because Transformer mechanism can only depend on relations of pixels to extract salient information [10], three fully connected layers in the Transformer mechanism depend on sizes of given patches, which will result in high complexities when given patches is big. To overcome this issue, we use three convolutional layers with sizes of fixed kernels rather than three fully connected layers to eliminate native effects of three fully connected layers in terms of the number of parameters for image denoising. Its effectiveness can be verified as follows. We assume that parameters of each fully-connected layer are $C\times p^{2}\times C\times p^{2}$, where $C$ is channel number, $p$ is size of given patches ($p=6$ in the GTEBlock). According to that, we see that parameters of each fully-connected layer are strong relation with $p$. However, parameters of each convolutional layer are $C\times C\times k\times k$ , where $k$ denotes kernel size ($k=3$). According to parameter computation method, we can see that the number of parameters is not affected by given patch sizes. Also, the number of parameters from each convolutional layer only takes 1/144 that of each fully connected layer, which can reduce computational costs. Besides, its performance does not have any varying via ID 5 of GTEBlock with fully connected layers (FCL) and ID 6 of GTEBlock as shown in TABLE I. We can see that GTEBlock with three fully connected layers has obtained same PSNR as GTEBlock with three convolutional layers on Urban100 [24] for image denoising with noise level of 15.
@@ IV-C Experimental results
We use quantitative and qualitative analysis to test denoising performance of our HWformer. Quantitative analysis includes synthetic and real noisy image denoising, comparisons of complexity and evaluation of denoising time. Qualitative analysis is used to evaluate visual effects of predicted denoising images. For synthetic noisy image denoising, we use Block-matching and 3D filtering (BM3D) [38], trainable nonlinear reaction diffusion (TNRD) [39], targeted image denoising (TID) [40], a denoising CNN (DnCNN) [7], image restoration CNN (IRCNN) [41], fast and flexible denoising network (FFDNet) [42], neural nearest neighbors networks (N3Net) [43], fractional optimal control network (FOCNet) [44], DRUNet [45], DAGL [46], robust deformed denoising CNN (RDDCNN) [47], cross Transformer denoising CNN (CTNet) [30], CSformer [48], image restoration using swin Transformer (SwinIR) [11], Restoration Transformer (Restormer) [49], image processing Transformer (IPT) [10] and encoder-decoder-based Transformer (EDT) [50] on Urban100 [24], BSD68 [31] and Set12 [7] for gray synthetic noisy image denoising, on CBSD68 [31], Kodak24 [33], SPARCS [34], McMaster [32] and Urban100 [24] for color synthetic noisy image denoising. To comprehensively validate the proposed method in image denoising, peak signal-to-noise ratio (PSNR)[51], structural similarity index (SSIM)[51], feature similarity index measure (FSIM)[52], learned perceptual image patch similarity (LPIPS) [53], peak signal-to-blur ratio (PSBR)[54] and color difference [55] are used as metric to conduct experiments. For gray image denoising, experiments of different methods for images denoising with single class and multiple classes are conducted. As shown in TABLE III, our HWformer method almost has obtained the best result for single image denoising with eleven classes when noise level is 15, 25 and 50, respectively. For instance, our HWformer exceeds 0.31dB than that of the second SwinIR for a Starfish image denoising with noise level of 50 in TABLE III. As shown in TABLE IV, we can see that our HWformer has nearly obtained the highest PSNR on BSD68 and Urban 100 for image denoising with multiple classes when noise level is 15, 25 and 50, respectively. For instance, our HWformer has an improvement of 0.15dB than that of the second Restormer on Urban 100 for noise level of 25 in TABLE IV. According to mentioned illustrations, it is known that our HWformer is effective for gray image denoising.
@@
For real noisy image denoising, we compare with popular methods, i.e., BM3D [38], TID [40], DnCNN [7], RIDNet [56], VDN [57], AINDNet [20], SADNet [58], RDDCNN [47], DANet [59], DeamNet [60], CycleISP [61], DAGL [46] and MPRNet [62] on CC [35] and SIDD [29]. Specifically, we use given reference images (Ground truth) from chosen real noisy image datasets and obtained clean images from different noisy models to compute PSNR values [51] to test denoising performance of our HWformer. As shown in TABLE X, our method has achieved the best performance besides one setting of Nikon D800 ISO = 1600 on CC for real noisy image denoising. For instance, our method has an improvement of 3.38dB than that of the second CBM3D in terms of PSNR values of the third image with the setting of Nikon D600 ISO = 3200 for real noisy image denoising in TABLE X. As TABLE XI, we can see that the proposed HWformer is competitive in contrast with other popular methods, i.e., ADGL and MPRNet on SIDD for real noisy image denoising.Also, red and blue lines denote the best and second results in TABLEs III-XI.
@@
Qualitative analysis: We choose an area of denoising images from different methods to amplify it as observation area. If observation area is clearer, its corresponding method is more effective for image denoising. Also, we choose DnCNN, SwinIR, Restormer, DRUNet, DGAL and our HWformer on BSD68 and Urban100 for noise levels of 50 to obtain gray clean images. As shown in Figs. 2 and 3, we can see that our HWformer has obtained clearer areas than that of other comparative methods. For comparisons of color visual images, we choose DnCNN, SwinIR, Restormer, DRUNet, EDT-B and our HWformer on McMaster and Urban100 for noise levels of 50 to obtain color clean images. As shown in Figs. 4 and 5, we can see that our HWformer can obtain clearer detailed information than that of other methods. For real image denoising, we can observe our HWformer is capable of restoring more details and texture information as shown in Fig. 6. Thus, our HWformer is effective for qualitative analysis.